OpenStudy (anonymous):

solve the equatio.. please show work so i can understand....2-3/4x=20

OpenStudy (anonymous):

2 - (3/4x) = 20 8x - 3 = 80x -3 = 72x x = -3/72

OpenStudy (anonymous):

-3/4 x = 20 - 2 = 18 -3x = 18*4 =72 x = -72/3 = -24 Check: 2 - 3/4 *-24 = 20 OK Assumes its three quarters x not 3 over 4x.